OpenStudy (anonymous):

determine the products [ 2 -3 5] [1 -1 1] [ 3 2-4] [1 -2 -2] [1 1-2 ] [2 1 3] and use it to solve the system of equations : x-y =4,x-2y-2z=9,2x+y+3z=1.

OpenStudy (anonymous):

How does the product help you solve for the systems? Just set up the augment matrix of the system of equations

OpenStudy (anonymous):

what is augment matrix? @Romero

OpenStudy (anonymous):

[1 -1 1 4 ] [1 -2 -2 9 ] [2 1 3 1 ]

OpenStudy (anonymous):

x-y+ z =4 x-2y-2z=9 2x+y+3z=1

OpenStudy (anonymous):

but i have never solved any matrix of 4 terms. so i can't use this one. :(

OpenStudy (anonymous):

Have you ever solved using row reduction?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Same concept is applied in larger matrices. It doesn't change....you just have to work a little longer.

OpenStudy (anonymous):

can you please solve this one.here please

OpenStudy (anonymous):

It's all the same [1 -1 1 4 ] [1 -2 -2 9 ] [2 1 3 1 ] [1 -1 1 4 ] [0 -1 -3 5 ] Here I subtracted the first row to row two and row three [0 3 1 -7] since the third had a 2 in it's first column I subtract 2*(row 1) [1 -1 1 4 ] [0 1 3 -5 ] I multiplied row two by -1 [0 3 1 -7] [1 0 4 -1 ] [0 1 3 -5 ] I subtract 3*(row 2) to row three [0 0 -8 8 ] and added row two to row 1

OpenStudy (anonymous):

Can you do the rest? Does this look familiar to you now?

OpenStudy (anonymous):

do i need to solve the equations now? @Romero

OpenStudy (anonymous):

What type of math are you taking?

OpenStudy (anonymous):

@Romero . i am using algebraic mathematics